# On the Complexity of Bipartite Degree Realizability

### v1

@@ 3 Polynomial algorithm for $\textsc{BDR}_{c_{1},c_{2}}$
where $a$ is positive. The linear function $f(z)$ is positive on the interval $0,\max\{z\}]$, if it is positive at the endpoints of the interval.(Recall that we assume that $n_{1}+m_{1}\geq n_{2}+m_{2}$, therefore $z\geq 0$.) We know that it is positive when $z=0$ for $a$ is positive. When $z$ takes its maximum, then the shorter degree sequence contains only $c_{2}n$ values and at most one $d_{2}$. However, we already have shown that in that case the number of $c_{2}n$ degrees is at least $c_{2}n$, therefore, the Gale-Ryser inequality is satisfied. This is equivalent with $f(z)$ is non-negative at the $\max\{z\}$ endpoint. It follows that the inequality in equation 6 always holds.
@@
$n_{1}c_{2}n+d_{1}\;\leq\;n_{2}\min\{n_{1}+1,c_{2}n\}\;+\;\min\{n_{1}+1,d_{2}\}\;+\;m_{2}\min\{n_{1}+1,c_{1}n\},$ ($\star\star$)
under the standing assumptions and with the notation of the theorem. Again, we prove ($\star\star$) by splitting into three exhaustive cases according to the relative size of $n_{1}$.
@@
Here $\min\{n_{1}+1,c_{2}n\}=c_{2}n$. Also $d_{2}<c_{2}n\leq n_{1}+1$, so $\min\{n_{1}+1,d_{2}\}=d_{2}$. Finally, $n_{1}+1\geq c_{2}n>c_{1}n$, hence $\min\{n_{1}+1,c_{1}n\}=c_{1}n$. Therefore the right-hand side of $(\star\star)$ equals
@@
Thus $(\star\star)$ holds in Case (A).
@@
which is equivalent to $c_{1}\geq\dfrac{c_{2}^{2}}{1-c_{2}}$ (the hypothesis of the theorem). Thus $(\star\star)$ holds in Case (B).
@@
$n_{1}c_{2}n+d_{1}\leq\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)(n_{1}+1)+\min\{n_{1}+1,d_{2}\}+\left(\frac{n}{2}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
@@
The expression $R_{1}$ in equation 3 is increased by a positive term ($\frac{\eta}{n}$) compared to the expression $R_{1}$ in equation 6, therefore it is positive. $\eta+\frac{1}{n}$ in expression $R_{2}$ in equation 3 plays the same role than $\eta$ in expression $R_{2}$ in equation 6, therefore $R_{2}$ is positive. That is, the right hand side is again a linear function $a+bz$ with $a>0$. Again, it is non-negative on the whole interval $0,\max\{z\}]$ if it is non-negative at the endpoints. It is positive when $z=0$ for $a>0$, and when $z$ has maximum, the shorter side of the bipartite degree sequence contains at least $c_{2}n$ number of $c_{2}n$ degrees, therefore, the Gale-Ryser inequality holds, therefore $a+bz$ is non-negative when $z$ takes its maximum. Therefore, the second critical Gale-Ryser inequality also holds, therefore, the degree sequence $D=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}nc_{2}n))$ is graphic.
We proved that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence $D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$ is graphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ is graphic if and only if it can be partitioned into two degree sequences with equal sum. Since it can be checked in polynomial time if a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ can be split into two degree sequences with equal sums, the $\textsc{BDS}_{c_{1},c_{2}}$ is in P for all $0<c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$. ∎
@@ 4 The conditional NP-completeness result
Now we turn to the other main result of the paper. Before we prove the main theorem, we state and prove a bunch of technical observations. They do not go beyond high school level mathematics, but on the other hand, they are not trivial, and we give formal proof for the sake of completeness.
Observation 2.
For $0<c_{1}<c_{2}<\frac{1}{2}$,
@@
Now we prove the reverse implication.
Assume
@@
We first prove the forward implication.
Suppose
@@
Now we prove the reverse implication.
Assume
@@
Proof.
Assume $0<c_{1}<c_{2}<\tfrac{1}{2}$.
We prove the equivalence.
First, suppose
@@
Theorem 3.
If BDR is in NP-complete, then $\textsc{BDR}_{c_{1},c_{2}}$ is also NP-complete for all $0<c_{2}<\frac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$.
Proof.
@@
First we need a technical computation that depends only on $c_{1}$ and $c_{2}$ and not on $D$ (therefore, it can be done in constant time). Recall that $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$ is equivalent with $c_{2}>\sqrt{2c_{1}}-c_{1}$, according to Observation 2. Let $c_{2}^{\prime}:=\sqrt{2c_{1}}-c_{1}$. Since $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$, it follows that $c_{2}^{\prime}<c_{2}$. Let $r$ be a rational number between $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ obtained in the following way: in the decimal representation of $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ let $k$ be the first digit in which the said two numbers differ. Take $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ rounded to $k$ digits (these are rational numbers) and take the arithmetic average of them. Let $r$ be the so obtained rational number. Then $c_{2}^{\prime}<\frac{1-r^{2}}{2}<c_{2}$. Since the function $1-x-\sqrt{1-2x}$ is strictly increasing on the interval $0,\frac{1}{2}]$, it follows that $c_{1}<\frac{(1-r)^{2}}{2}\left(=1-\frac{1-r^{2}}{1}-\sqrt{1-2\frac{1-r^{2}}{2}}\right)<1-c_{2}-\sqrt{1-2c_{2}}$. For short, we define $\tilde{c_{1}}:=\frac{(1-r)^{2}}{2}$ and $\tilde{c_{2}}:=\frac{1-r^{2}}{2}$. Since both $\tilde{c_{1}}$ and $\tilde{c_{2}}$ are rational, for infinitely many $n$, $\tilde{c_{1}}n$, $\tilde{c_{2}}n$ and $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ are all integers. Furthermore, those $n$’s are a multiple of a given integer, depending on $r$.
@@
$\frac{\tilde{c_{2}}n+\frac{S}{2}}{n+S}<c_{2}$ (8)
@@
$\frac{\tilde{c_{1}}n}{n+S}>c_{1}.$ (9)
Such an $n$ exists since $\tilde{c_{1}}>c_{1}$ and $\tilde{c_{2}}<c_{2}$. Further, this $n$ is a linear function of $S$ (and thus, a polynomial function of $|D|$). Indeed, the inequalities in equations 8 and 9 holds for
@@
and we require that $n$ be a multiple of an integer depending on $r$ and $S$. Then the degree sequence $D^{\prime}$ contains $(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$, for each $d_{i}\in D$, a degree $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$, S|D|S-|D| instances of degree $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ and $n-(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{1}}n$. That is, $D^{\prime}$ has length $n+S$, and due to equations 8 and 9, $D^{\prime}$ is a problem instance in $\textsc{BDR}_{c_{1},c_{2}}$. Further, it can be constructed from $D$ in polynomial time.
@@
Indeed, this is equivalent with $\tilde{c_{1}}=1-\tilde{c_{2}}-\sqrt{1-2\tilde{c_{2}}}$, according to Observation 3. Similarly, create a semi-regular graph between $V_{L}$ and $U_{S}$ with equivalent degrees. The so-obtained graph is a bipartite graph realization of $D^{\prime}$. Indeed, the degrees in $U_{L}$ (in $V_{L}$, respectively) are the sum of the degrees obtained from the two complete bipartite graphs towards $V_{L}$ and $V$ (to $U_{L}$ and $U$, respectively) and the degrees towards $U_{S}$ (respectively, $V_{S}$):
@@
$\frac{(\tilde{c_{1}}+\tilde{c_{2}})n\left(\tilde{c_{2}}n+\frac{S}{2}\right)+(1-\tilde{c_{1}}-\tilde{c_{2}})n\tilde{c_{1}}n+S\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+S}{2}-\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\tilde{c_{2}}n+\frac{S}{2}\right)=$
@@
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t+x\right)+t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}=$
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\right)^{2}-t^{2}+x\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)+t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}$
@@
which is positive. Therefore in any realization $G=(U,V,E)$ of $D^{\prime}$, there are $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ instances of degree $\tilde{c_{2}}+\frac{S}{2}$ both in $U$ and in $V$. It follows that the left-hand side and the right-hand side are the same in the $k^{\mathrm{th}}$ Gale-Ryser inequality with $k=\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$. Further, it follows, if we remove the edges incident to vertices having degree $\tilde{c_{2}}+\frac{S}{2}$, the remaining graph will be a realization of $D$ amalgamated with several further vertices having degree 0. Therefore, $D$ has a bipartite graph realization. ∎
Acknowledgement
The work was supported by the Hungarian NKFIH grant K132696 and by the European Union project RRF2.3.1-21-2022-00006 within the framework of Health Safety National Laboratory Grant no RRF-2.3.1-21-2022-00006.

### v2

@@ 3 Polynomial algorithm for $\textsc{BDR}_{c_{1},c_{2}}$
where $a$ is positive. The linear function $f(z)$ is positive on the interval $0,\max\{z\}]$, if it is positive at the endpoints of the interval.(Recall that we assume that $n_{1}+m_{1}\geq n_{2}+m_{2}$, therefore $z\geq 0$.) We know that it is positive when $z=0$ for $a$ is positive. When $z$ takes its maximum, then the shorter degree sequence contains only $c_{2}n$ values and at most one $d_{2}$. However, we already have shown that in that case the number of $c_{2}n$ degrees is at least $c_{2}n$, therefore, the Gale-Ryser inequality is satisfied. This is equivalent with $f(z)$ is non-negative at the $\max\{z\}$ endpoint. It follows that the inequality in equation 3.3 always holds.
@@
(9) $n_{1}c_{2}n+d_{1}\;\leq\;n_{2}\min\{n_{1}+1,c_{2}n\}\;+\;\min\{n_{1}+1,d_{2}\}\;+\;m_{2}\min\{n_{1}+1,c_{1}n\},$
under the standing assumptions and with the notation of the theorem. Again, we prove equation 9 by splitting into three exhaustive cases according to the relative size of $n_{1}$.
@@
Here $\min\{n_{1}+1,c_{2}n\}=c_{2}n$. Also $d_{2}<c_{2}n\leq n_{1}+1$, so $\min\{n_{1}+1,d_{2}\}=d_{2}$. Finally, $n_{1}+1\geq c_{2}n>c_{1}n$, hence $\min\{n_{1}+1,c_{1}n\}=c_{1}n$. Therefore the right-hand side of equation 9 equals
@@
Thus equation 9 holds in Case (A).
@@
which is equivalent to $c_{1}\geq\dfrac{c_{2}^{2}}{1-c_{2}}$ (the hypothesis of the theorem). Thus equation 9 holds in Case (B).
@@
$n_{1}c_{2}n+d_{1}$ $\leq$ $\left(n_{1}+\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}\right)(n_{1}+1)+$
$+\min\{n_{1}+1,d_{2}\}+\left(\frac{n}{2}-x-\frac{\frac{d_{1}-d_{2}}{n}+2xc_{1}}{c_{2}-c_{1}}-1\right)c_{1}n.$
@@
The expression $R_{1}$ in equation 3.3 is increased by a positive term ($\frac{\eta}{n}$) compared to the expression $R_{1}$ in equation 3.3, therefore it is positive. $\eta+\frac{1}{n}$ in expression $R_{2}$ in equation 3.3 plays the same role than $\eta$ in expression $R_{2}$ in equation 3.3, therefore $R_{2}$ is positive. That is, the right hand side is again a linear function $a+bz$ with $a>0$. Again, it is non-negative on the whole interval $0,\max\{z\}]$ if it is non-negative at the endpoints. It is positive when $z=0$ for $a>0$, and when $z$ has maximum, the shorter side of the bipartite degree sequence contains at least $c_{2}n$ number of $c_{2}n$ degrees, therefore, the Gale-Ryser inequality holds, therefore $a+bz$ is non-negative when $z$ takes its maximum. Therefore, the second critical Gale-Ryser inequality also holds, therefore, the degree sequence $D=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}nc_{2}n))$ is graphic.
We proved that for any $s_{1}\leq s_{2}$ with $s_{1}+s_{2}=n$, and for any $\Sigma$ with $c_{1}s_{2}\leq\Sigma\leq c_{2}s_{1}$, the bipartite degree sequence
$D:=(LBDS(s_{1},\Sigma,c_{1}n,c_{2}n),LBDS(s_{2},\Sigma,c_{1}n,c_{2}n))$
is bigraphic. Therefore, a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ has a bipartite graph realization if and only if it can be partitioned into two degree sequences with equal sum. Since it can be checked in polynomial time if a degree sequence $D\in\mathcal{D}(n,2\Sigma,c_{1}n,c_{2}n)$ can be split into two degree sequences with equal sums, the $\textsc{BDS}_{c_{1},c_{2}}$ is in P for all $0<c_{1}\leq c_{2}\leq\frac{\sqrt{c_{1}(c_{1}+4)}-c_{1}}{2}$.
@@ 4 The conditional NP-completeness result
Now we turn to the other main result of the paper. Before we prove the main theorem, we state and prove a bunch of technical observations. They do not go beyond high school level mathematics, but on the other hand, they are not trivial, and we give formal proof for the sake of completeness. {observation} For $0<c_{1}<c_{2}<\frac{1}{2}$,
@@
Now we prove the reverse implication. Assume
@@
We first prove the forward implication. Suppose
@@
Now we prove the reverse implication. Assume
@@
Proof 4.3.
Assume $0<c_{1}<c_{2}<\tfrac{1}{2}$. We prove the equivalence. First, suppose
@@
Theorem 4.4.
If BDR is NP-complete, then $\textsc{BDR}_{c_{1},c_{2}}$ is also NP-complete for all $0<c_{2}<\frac{1}{2}$ and $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$.
Proof 4.5.
@@
First we need a technical computation that depends only on $c_{1}$ and $c_{2}$ and not on $D$ (therefore, it can be done in constant time). Recall that $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$ is equivalent with $c_{2}>\sqrt{2c_{1}}-c_{1}$, according to Observation 4. Let $c_{2}^{\prime}:=\sqrt{2c_{1}}-c_{1}$. Since $c_{1}<1-c_{2}-\sqrt{1-2c_{2}}$, it follows that $c_{2}^{\prime}<c_{2}$. Let $r$ be a rational number between $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ obtained in the following way: in the decimal representation of $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ let $k$ be the first digit in which the said two numbers differ. Take $\sqrt{1-2c_{2}}$ and $\sqrt{1-2c_{2}^{\prime}}$ rounded to $k$ digits (these are rational numbers) and take the arithmetic average of them. Let $r$ be the so obtained rational number. Then $c_{2}^{\prime}<\frac{1-r^{2}}{2}<c_{2}$. Since the function $1-x-\sqrt{1-2x}$ is strictly increasing on the interval $0,\frac{1}{2}]$, it follows that $c_{1}<\frac{(1-r)^{2}}{2}\left(=1-\frac{1-r^{2}}{1}-\sqrt{1-2\frac{1-r^{2}}{2}}\right)<1-c_{2}-\sqrt{1-2c_{2}}$. For short, we define $\tilde{c_{1}}:=\frac{(1-r)^{2}}{2}$ and $\tilde{c_{2}}:=\frac{1-r^{2}}{2}$. Since both $\tilde{c_{1}}$ and $\tilde{c_{2}}$ are rational, for infinitely many $n$, $\tilde{c_{1}}n$, $\tilde{c_{2}}n$ and $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ are all integers. Furthermore, those $n$’s are a multiple of a given integer, depending on $r$.
@@
(11) $\frac{\tilde{c_{2}}n+\frac{S}{2}}{n+S}<c_{2}$
@@
(12) $\frac{\tilde{c_{1}}n}{n+S}>c_{1}.$
Such an $n$ exists since $\tilde{c_{1}}>c_{1}$ and $\tilde{c_{2}}<c_{2}$. Further, this $n$ is a linear function of $S$ (and thus, a polynomial function of $|D|$). Indeed, the inequalities in equations 11 and 12 holds for
@@
and we require that $n$ be a multiple of an integer depending on $r$ and $S$. Then the degree sequence $D^{\prime}$ contains $(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{2}}n+\frac{S}{2}$, for each $d_{i}\in D$, a degree $d_{i}+\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$, S|D|S-|D| instances of degree $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ and $n-(\tilde{c_{1}}+\tilde{c_{2}})n$ instances of degree $\tilde{c_{1}}n$. That is, $D^{\prime}$ has length $n+S$, and due to equations 11 and 12, $D^{\prime}$ is a problem instance in $\textsc{BDR}_{c_{1},c_{2}}$. Further, it can be constructed from $D$ in polynomial time.
@@
Indeed, this is equivalent with $\tilde{c_{1}}=1-\tilde{c_{2}}-\sqrt{1-2\tilde{c_{2}}}$, according to Observation 4. Similarly, create a semi-regular graph between $V_{L}$ and $U_{S}$ with equivalent degrees. The so-obtained graph is a bipartite graph realization of $D^{\prime}$. Indeed, the degrees in $U_{L}$ (in $V_{L}$, respectively) are the sum of the degrees obtained from the two complete bipartite graphs towards $V_{L}$ and $V$ (to $U_{L}$ and $U$, respectively) and the degrees towards $U_{S}$ (respectively, $V_{S}$):
@@
$\frac{(\tilde{c_{1}}+\tilde{c_{2}})n\left(\tilde{c_{2}}n+\frac{S}{2}\right)+(1-\tilde{c_{1}}-\tilde{c_{2}})n\tilde{c_{1}}n+S\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+S}{2}$-
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\tilde{c_{2}}n+\frac{S}{2}\right)=t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}{n}+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}$
@@
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+t+x\right)+$
$t\left(\tilde{c_{2}}n+\frac{S}{2}\right)+\frac{1-\tilde{c_{1}}-\tilde{c_{2}}}{2}n\tilde{c_{1}}n+\frac{S}{2}\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n+\frac{S}{2}=$
$\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n\right)^{2}-t^{2}+x\left(\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n-t\right)+$
@@
which is positive. Therefore in any realization $G=(U,V,E)$ of $D^{\prime}$, there are $\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$ instances of degree $\tilde{c_{2}}+\frac{S}{2}$ both in $U$ and in $V$. It follows that the left-hand side and the right-hand side are the same in the $k^{\mathrm{th}}$ Gale-Ryser inequality with $k=\frac{\tilde{c_{1}}+\tilde{c_{2}}}{2}n$. Further, it follows, if we remove the edges incident to vertices having degree $\tilde{c_{2}}+\frac{S}{2}$, the remaining graph will be a realization of $D$ amalgamated with several further vertices having degree 0. Therefore, $D$ has a bipartite graph realization.